# Inter-turbine Modelling of Wind-Farm Power using Multi-task Learning

### v1

@@ 7 Conclusions
A probabilistic, multivariate wind-turbine power-prediction model was developed, utilising a beta regression model with B-spline basis functions (named the BSBR model). This BSBR model was shown to be successful in capturing both the heteroscedastic and asymmetric properties of power-curve data. When training the BSBR model on each turbine within a wind-farm, correlations were observed in the learned model parameters with respect to turbine locations, which were deemed to be associated with the directional dependency of turbine wakes.
@@
These spatial correlations in parameters were then leveraged via a “metamodel”, which imposed functional constraints over turbine-specific parameters, as part of a partially-pooled hierarchical Bayesian model. This model design is able to learn from all turbines (tasks) simultaneously, and via the metamodel parameters, can be used to make adaptive probabilistic predictions for turbines not necessarily included in the training data, using their spatial coordinates alone. The metamodel was found to outperform three benchmark models that utilised no-pooling, partial-pooling and complete-pooling of information, both in terms of normalised mean squared error, and joint log-likelihood.
@@
While for power-curve prediction, training data are readily available for all turbines, the results suggest that the metamodel design is an efficient modelling solution for situations where data may be limited, and correlations could be expected in a variable of interest among a population of structures. Potential further applications in this domain include (but are not limited to): structural load assessment for fatigue-life estimation, condition monitoring and fault detection, and design for wind-farm layouts.
@@
A number of simplifying assumptions were made in relation to the training features to help capture wake effects from the data; these features may be improved if meteorological mast data were available, which would likely be a more accurate reflection of freestream wind conditions. Given the complexity of turbine wake physics, future work could also consider using physics-based inputs, to develop a hybrid data and physics based approach. Datasets labelled with turbine operating regimes, and instances of damage or repair that may indicate a change in underlying turbine behaviour, may also help with data preprocessing to reduce possible sources of noise. Model design may also be improved via optimisation of the number of B-spline knots per feature, their locations, or alternative spline-modelling approaches. The use of nonlinear metamodels may also worth considering, particularly if parameter spatial correlations are complex.

### v2

@@ 7 Conclusions
A probabilistic, multivariate wind-turbine power-prediction model was developed, utilising a beta regression model with B-spline basis-functions (named the BSBR model). This BSBR model was shown to be successful in capturing both the heteroscedastic and asymmetric properties of power-curve data. When training the BSBR model on each turbine within a wind-farm, correlations were observed in the learned model parameters with respect to turbine locations, which were deemed to be associated with the directional dependency of wind-turbine wakes.
@@
These spatial correlations in parameters were then leveraged via a “metamodel”, which imposed functional constraints over turbine-specific parameters, as part of a hierarchical Bayesian model. This model design is able to learn from all turbines (tasks) simultaneously, and via the metamodel parameters, can be used to make adaptive probabilistic predictions for turbines not necessarily included in the training data, using their spatial coordinates alone. Across an 80-turbine wind-farm, the metamodel was found to outperform three benchmark models that utilised no-pooling, partial-pooling and complete-pooling of information, both in terms of normalised mean squared error (NMSE) and joint log-likelihood (JLL) in the majority of cases, while having a relatively low number of model parameters.
Since the models are trained on farm-level input features to predict turbine-level output power, they are tuned (to a degree) on the specific environmental conditions (including wake-effects) of the wind-farm used in this work, which are also related to the specific layout of the wind-farm, as well as the turbine-type. If one wishes to apply this model to another wind-farm of interest, the model parameters should therefore be re-learned on data directly from that wind-farm in order to obtain valid results.
@@
While in practice, training data for all turbines in a wind-farm are often available for power-curve prediction, the results suggest that the metamodel design is an efficient modelling solution for situations where data may be limited, and correlations could be expected in a variable of interest among a population of structures. Potential further applications in this domain include (but are not limited to): structural load assessment for fatigue-life estimation, condition monitoring and fault detection, and design for wind-farm layouts. As another perspective, suppose one wishes to measure a new variable of interest in a population of structures; this modelling approach could justify instrumenting just a subset of this population. By leveraging shared information across the subset of the population, the model may estimate the variable of interest with sufficient precision for the remaining structures without sensors, reducing the overall cost of sensor deployment.
As highlighted in Section 6.1, the complexity of wake-effects cannot be fully captured by the wind-direction alone, which likely limits the performance of all the models in predicting power in the region between the cut-in and the rated wind-speeds. Future work could consider incorporating additional explanatory variables, such as wind-shear, turbulence-intensity, thermal-stratification and air-density variations, which may improve predictive performance. For the turbulence-intensity, if the standard-deviation of the wind-speed was available in the SCADA data (which is sometimes the case), it may be estimated as the ratio of the standard-deviation to the mean wind-speed [38]. The incorporation of weather forecast (or hindcast) data may also help in gaining information for the additional variables, which could be obtained from a provider such as ECMWF [25]. The additional uncertainty associated with forecasted data does need to be considered however, as well as how to combine the 10-minute SCADA data with the weather data, which is likely to be coarser. If turbine-specific data are used, then the incorporation of computational fluid dynamics-based wake modelling—or a surrogate model with lower computational cost—may offer further improvements. The PyWake wind-farm simulation tool [39] may be a practical option for such surrogate modelling.
@@
A number of simplifying assumptions were made in relation to the training features, to help capture wake-effects from the data; these features may be improved if meteorological mast data were available, which would likely be a more accurate reflection of freestream wind conditions. Given the complexity of turbine wake physics, future work could also consider using physics-based inputs, to develop a hybrid data and physics-based approach. Datasets labelled with turbine operating regimes, and instances of damage or repair that may indicate a change in underlying turbine behaviour, may also help with data preprocessing to reduce possible sources of noise. Model design may also be improved by optimising the number of B-spline knots per feature, or their locations—perhaps by increasing the density of knots in the regions of the cut-in wind-speed and where rated power is reached, since these regions have increased curvature. Alternative spline-modelling approaches could also be considered. The use of nonlinear metamodels may also be worth considering, particularly if parameter spatial correlations are complex.
If one wishes to scale the model for larger wind-farms, or perhaps with more data, it may be practical to reduce the computational cost. One approach could be the use of variational inference (instead of HMC), where the posterior is approximated as one of a family of known distributions [31]; this however may come at the cost of accuracy. Keeping the total number of model parameters as low as reasonably practicable is also likely to reduce the computational burden.